OpenStudy (anonymous):

how do I rewrite ln x=7 into an exponential equation?

OpenStudy (campbell_st):

well use the log law \[\log_a(x) = b\] then raising each side of the equation as a power of the base \[a^{\log_{a}(x)} = a^b\] which simplifies to \[x = a^b\]

OpenStudy (anonymous):

thanks!